2. Comparative Advantage

Continue using the information about Timothy and John from the previous question. ______________ has the comparative advantage in producing shoes and _____________ has the comparative advantage in producing clothes.

A) None of the other answers are correct

B) John, Timothy

C) John, John

D) Timothy, Timothy

E) Timothy, John
Answer #1

E) Timothy, John

Explanation:

For Timothy, the opportunity cost of 1 pair of shoes = 4/2 = 2 pair of clothes.

For Timothy, the opportunity cost of 1 pair of clothes = 2/4 = 0.5 pair of shoes.

For John, the opportunity cost of 1 pair of shoes = 5 pair of clothes.

For John, the opportunity cost of 1 pair of clothes = 1/5 = 0.2 pair of shoes.

So, Timothy has a lower opportunity cost in the production of shoes i..e Timothy has a comparative advantage in shoes.

So, John has a lower opportunity cost in the production of clothes i..e John has a comparative advantage in clothes.